HIGH RESOLUTION NEAR INFRARED SPECTRA OF THE CH STRETCHING MODES IN JET COOLED ETHYL RADICAL

\underline{SCOTT DAVIS}, DAIRENE UY and DAVID J. NESBITT, JILA, National Institute of Standards and Technology and Department of Chemistry and Biochemistry, University of Colorado, Boulder, Colorado 80309-0440.